# What Does Neuro Mean to Cardio? Investigating the Role of Clinical Specialty Data in Medical LLMs

### v2

@@ Abstract
In this paper, we introduce S-MedQA, an English medical question-answering (QA) dataset for benchmarking large language models in fine-grained clinical specialties. We use S-MedQA to check the applicability of a popular hypothesis related to knowledge injection in the knowledge-intense scenario of medical QA, and show that 1) training on data from a specialty does not necessarily lead to best performance on that specialty and 2) regardless of the specialty fine-tuned on, token probabilities of clinically relevant terms for all specialties increase consistently. Thus, we believe improvement gains come mostly from domain shifting (e.g., general to medical) rather than knowledge injection and suggest rethinking the role of fine-tuning data in the medical domain. We release S-MedQA and all code needed to reproduce all our experiments to the research community.  [^1] https://github.com/yanxinlan/S-MedQA
@@ 1 Introduction
Multiple-choice question-answering (QA) datasets are widely used to benchmark large language models (LLMs) in the medical domain (Singhal et al., 2023; Labrak et al., 2024) and guide the development of medical LLMs (e.g., PubMedQA, Jin et al., 2019; MedQA, Jin et al., 2021; MedMCQA, Pal et al., 2022). However, specialized hospitals may require LLMs to address specific clinical problems and are often interested in performance within one or a few clinical specialties (e.g., obstetrics or oncology). To the best of our knowledge, no open-source medical QA datasets include clinical specialty annotations, limiting research into knowledge transfer across clinical specialties.
To address this gap, we develop S-MedQA, the first English medical QA dataset with multiple clinical specialty annotations. We build S-MedQA based on the widely used MedQA (Jin et al., 2021) and MedMCQA (Pal et al., 2022) datasets, and use gpt-3.5-turbo-0125 (GPT-3.5) and medical experts to map samples onto clinical specialties. We label QA pairs with single specialties using tailored prompts, retaining only majority-agreed annotations. Expert validation on single specialty achieves up to 97.8% accuracy (details in §2.4). S-MedQA spans 15 specialties, each with hundreds to thousands of samples (see §2 for details). We then expand these annotations to multiple specialties using conformal prediction with calibrated confidence guarantees, achieving a coverage rate of 95% and an F-1 score of 0.60 at a 95% confidence level on a curated calibration set.
We use our dataset to investigate the hypothesis that almost all knowledge in LLMs originates from pretraining, and that fine-tuning primarily serves to shift the model toward a specific knowledge domain (Zhou et al., 2024). To that regard, we first fine-tune LLMs on one specialty and evaluate on other specialties (see §3.2). Interestingly, the best results often come from fine-tuning on unrelated specialties. E.g., fine-tuning with infectious disease data performs best on cardiology domain, despite their knowledge being largely unrelated. Moreover, different pre-trained LLMs exhibits different knowledge transfer patterns across clinical specialties (see §A.12). In further experiments, we curate clinically relevant per-specialty terms (e.g., diseases, procedures) and analyze changes in these term probabilities before and after fine-tuning on data from different specialties. Our results suggest that performance gains are driven more by domain shifts (from general to medical domain) than by fine-grained clinical knowledge injection.
@@
Figure 1: Overview of S-MedQA’s construction process. For single specialty annotation of each sample, we generate predictions using 5 different prompts and only keep those where predictions agree ($3+$, $4+$, or $5$ times). We randomly sample $1,000$ questions from our train set and ask a medical expert to evaluate GPT-3.5’s predictions, achieving an accuracy ranging from $97.8$% to $90.8$%. Three medical students then annotate the same $200$ samples out of the original $1,000$ annotated samples for computing inter-annotator agreements (see §2.4 for details). The expert also manually annotates S-MedQA’s whole validation and test set. For multi-specialty annotation, we leverage conformal prediction to assign labels across multiple clinical specialties, using a calibration set of $300$ samples manually annotated by the medical expert and medical students. At a target confidence level of 90%, we determine a conformal threshold of $0.5739$, which is then applied to all remaining samples for calibrated predictions.
2 A Benchmark of clinical specialties
@@
We describe the creation of S-MedQA, a high-quality benchmark for medical QA with clinical specialty annotations (overview in Fig. 1). We release multiple versions of S-MedQA with varying accuracy/coverage trade-offs, controlled by majority voting thresholds for including examples. Users can opt for a cleaner dataset with fewer samples or a noisier one with more samples (details in §2.4).
Data and splits
@@ 2.1 Clinical specialty categorization
Single clinical specialties
Manually labeling QA examples with clinical specialties is costly and time-consuming. To reduce manual effort, we use GPT-3.5 to annotate samples with single clinical specialties. However, preliminary experiments with a single prompt to predict a single specialty achieve low accuracy($\sim 75$%). To improve this, we design five prompts (details in §A.2), generate single specialty predictions with GPT-3.5 for each Q&A sample, and apply majority voting to select the specialty (Ding et al., 2023; Goel et al., 2023). Human evaluation shows accuracies between $90.8$%–$97.8$% (see §2.4 for details).
From single to multiple clinical specialties
Single specialty labels do not always capture the complexity of clinical questions. To account for this, we applied conformal prediction (Angelopoulos and Bates, 2021) to generate calibrated multi-specialty labels in S-MedQA. At a 95% confidence level, we achieve an average coverage of 95% and an F1 score of 0.60 on the calibration set (see §A.5 for the methods and evaluation).
2.2 S-MedQA’s predicted specialties
In §A.3 (Fig. 3) we show the distribution of predicted specialties. We exclude $2310$ ($6$%) samples categorized as Others (containing clinically irrelevant information), and then focus on $15$ out of $55$ specialties with more than $300$ samples to ensure statistical reliability. The final dataset comprises $40007$ / $899$ / $893$ samples in train/validation/test sets after human validation (described next in §2.4). For subsequent analyses, we use the top 6 specialties with the most samples (Cardiology, Gastroenterology, Infectious diseases, Neurology, Obstetrics and Gynecology, and Pediatrics).
2.3 Term overlap analysis
We analyze the overlap of clinically relevant terms in the train/test splits of specialties (within and across specialties). We use scispaCy [^2] https://allenai.github.io/scispacy to extract medical terms from S-MedQA based on the Unified Medical Language System (Bodenreider, 2004). We map each term to relevant clinical specialties by automatically linking the medical concepts to disorders via the Human Phenotype Ontology (HPO; Castellanos et al., 2024) and tracing diseases to their ancestors in the SNOMED-CT [^3] https://www.snomed.org/ hierarchy to identify top-level systematic disease categories. General terms unrelated to specific specialties or shared by four or more of the six specialties are excluded from this analysis. Please refer to §A.8 for more details. In summary, the average overlap of clinically relevant terms within specialties is 63.4%, whereas across specialties is 32.8%, which means that term mentions are consistent within a specialty’s train/test splits, whereas different specialties share limited terminology.
@@ 2.4 Manual validation for single specialty predictions
A medical expert labels all the examples in S-MedQA’s validation and test sets with the single most correct clinical specialty. This expert also validates $1,000$ random samples from the train set, confirming whether the specialties predicted by GPT-3.5 are correct. [^4] This is the specialty that is the most relevant to the QA pair. In this evaluation, we exclude any cases where there is ambiguity in which clinical specialty is the most relevant. In general, when using voting with multiple prompts, we see large performance gains compared to using single prompts (e.g., from 72.872.872.872.880.280.280.280.2% to 90.890.890.890.897.897.897.897.8%; see §A.7).
# votes (out of 5)
$3+$ $4+$ $5$
@@
(Sample size) (80) (83) (102) (74) (88) (90)
Mistral-v0.2 52.0 45.9 48.2 37.0 52.9 43.5 46.9
Train Sets Cardio 51.8 54.6 44.3 47.6 51.5 44.6 49.4
Gastro 54.9 54.1 38.9 43.5 51.7 46.4 48.8
Infect 55.4 53.9 43.0 47.8 54.3 43.5 50.1
Neuro 54.4 54.8 41.9 45.7 56.0 49.0 50.8
Obstetrics 54.4 52.4 45.7 41.9 51.5 44.6 48.8
Pediatrics 50.8 53.0 42.4 35.9 48.1 46.1 46.5
Combined 53.7 53.8 42.7 43.7 52.2 45.7 49.1
Table 2: Accuracy matrix for Mistral-v0.2 as the base model. Model is applied without finetuning. Model is trained on the combination of all 6 specialty train sets. For each specialty, the best performance when fine-tuned on different specialty datasets is in bold, and scores for models fine-tuned on the same specialty are underlined. Surprisingly, none of the best performances come from models fine-tuned on their corresponding training sets.
In Table 1, we show the accuracy vs. coverage trade-off over the $1,000$ random samples from the train set for different requirements for majority voting. A higher quorum results in higher accuracy ($90.8\rightarrow 97.8$) but greatly decreases the coverage ($89.1\rightarrow 49.2$). We release per-prompt single-specialty categorizations together with votes for all examples for users of the dataset to decide their preference between accuracy and coverage—more data but possibly more noise or less noise but less data—based on their specific use cases. We select ‘3+’ as the quorum in this study for adequate fine-tuning data.
3 Cross-specialty evaluation
@@ 3.1 Experimental setup
We experiment on 8 open-source LLMs. Six general-purpose models: Llama2-Chat-7B, Llama2 -Chat-13B, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct (Touvron et al., 2023), Mistral-Instruct-v0.1 and v0.2 (Jiang et al., 2023), one biomedical model: Bio-Medical-LLaMA-3-8B. We also include OLMo-2-1124-7B-Instruct to rule out the possible contamination of training data.
We fine-tune each LLM on the six per-specialty training sets using prompts from §A.9, and evaluate on all six test sets. Models are trained for up to 10 epochs, with selection based on per-specialty validation accuracy. We also train on the combined dataset to evaluate how exposure to larger and more diverse data affects models’ performance. Additional details and hyperparameters are in §A.10.
@@
We follow best practices to evaluate LLM performance on QA datasets by shuffling answers multiple times and adding multiple shuffled QA pairs in the test set, and by using the entire answer to a question instead of the LLM’s maximum token probability among options A, B, C, D in the answer (Zheng et al., 2023; Wang et al., 2024). Please see §A.11 for details on our evaluation protocol. To verify the effectiveness of the fine-tuning process, we assess the model’s performance on the training data and achieve an average accuracy of 89.3% across the fine-tuning datasets.
3.2 Results
Table 2 shows the performance of Mistral-v0.2 fine-tuned independently on each specialty training set and tested on all six specialty test sets. We also report the performance after fine-tuning on the combined training set. We observe that the models fine-tuned on the combined dataset, as well as each single specialty, consistently outperform the base model in terms of average performance, demonstrating the effectiveness of instruction fine-tuning.
Are improvements in Table 2 truly indicative of knowledge acquisition or injection?

### v3

@@ Abstract
In this paper, we introduce S-MedQA, an English medical question-answering (QA) dataset for benchmarking large language models (LLMs) in fine-grained clinical specialties. S-MedQA has over 2020k examples, covers $15$ medical specialties, and QA pairs can have multiple specialty annotations (e.g., when a question is cross-disciplinary), constructed with both machine and expert verification to maximize data availability. We use S-MedQA to investigate the role of clinical specialty data in the knowledge-intensive scenario of medical QA. Our results show that 1) training on data from a clinical specialty does not necessarily lead to best performance on that specialty, and 2) regardless of the specialty the LLM was fine-tuned on, token probabilities of clinically relevant terms increase consistently across all specialties. Thus, we hypothesize improvement gains are derived mostly from domain shifting (e.g., general to medical) rather than specialty-specific knowledge injection, and suggest rethinking the role of fine-tuning data in the medical domain. To motivate further advancements in the clinical NLP field, we release S-MedQA and all code needed to reproduce all our experiments to the research community. [^1] https://anonymous.4open.science/r/S-MedQA-85FD
@@ 1 Introduction
Multiple-choice question-answering (QA) datasets are widely used to benchmark large language models (LLMs) in the medical domain (Singhal et al., 2023; Labrak et al., 2024) and guide the development of medical LLMs (e.g., PubMedQA, Jin et al., 2019; MASH-QA, Zhu et al., 2020; MedQA, Jin et al., 2021; MedMCQA, Pal et al., 2022). However, specialized hospitals may require LLMs to address specific clinical problems and are often interested in performance within one or a few clinical specialties (e.g., obstetrics or oncology). To the best of our knowledge, no open-source medical QA datasets include clinical specialty annotations.
To address this gap, we propose S-MedQA, the first English medical QA dataset with multiple clinical specialty annotations (see Figure 1 for an overview, and §2 for details). We build S-MedQA based on the widely used MedQA (Jin et al., 2021) and MedMCQA (Pal et al., 2022) datasets, incorporating both machine and expert verification to semi-automatically map samples onto clinical specialties at scale, while guaranteeing the quality of the annotations. We first use multiple prompts and a majority voting mechanism to label QA pairs with a single specialty. S-MedQA includes 15 specialties, each with hundreds to thousands of samples, and single specialty expert validation shows up to 97.8% accuracy in this annotation (see §2 for details). In order to account for cross-disciplinary questions, the second step in our data collection is to expand single-specialty annotations into multiple specialties. To do that, we use a multi-label conformal prediction procedure with coverage guarantees. We annotate a held-out set of QA pairs with medical experts and build a conformal multi-label annotation procedure that leads to a precision of $0.69$, recall of $0.52$, 2424% exactly correct predictions, and an average number of labels of $1.55$.
@@
Figure 1: Overview of S-MedQA’s construction process. For single specialty annotation of each sample, we generate predictions using 5 different prompts and only keep those where predictions agree ($3+$, $4+$, or 55 times). A medical expert manually annotates S-MedQA’s entire validation and test set. We randomly sample $1,000$ questions from our train set and ask multiple medical experts to evaluate GPT-3.5’s predictions (we detail inter-annotator agreement in §2.2), achieving accuracies ranging from 90.890.897.897.8%. For multi-specialty annotation, we leverage conformal prediction to assign labels across multiple clinical specialties, using a held-out set of $300$ samples manually annotated by medical experts as calibration/test set. It achieves a precision of $0.69$, recall of $0.52$, with $24\%$ exactly correct matches and an average prediction length of $1.55$.
Different clinical specialties can have very different amounts of available data (e.g., see Figure 2). We thus first use our dataset to address the following research question (§3.2): to what extent can LLMs use the knowledge learned from a clinical specialty to answer questions about other specialties? We investigate Zhou et al. (2024)’s hypothesis that almost all knowledge in LLMs originates from pretraining, and that fine-tuning primarily serves to shift the model toward a specific knowledge domain. To that regard, we first fine-tune LLMs on one clinical specialty and evaluate on all other specialties. Interestingly, we find that the best results often come from fine-tuning on unrelated clinical specialties. E.g., fine-tuning with neurology data performs best on cardiology domain, despite their knowledge being largely unrelated. Moreover, although different pre-trained LLMs exhibit different knowledge transfer patterns across clinical specialties, neither of the LLMs we investigated show the best knowledge transfer when training and testing within a same clinical specialty.
These results lead us to take a step back and ask: but are QA pairs from different clinical specialties really different? To answer this question, in §3.3 we curate the relevant clinical terms (e.g., diseases, procedures) for each clinical specialty using authoritative resources—such as the SNOMED-CT knowledge graph Cornet and de Keizer (2008)—to effectively estimate the clinical knowledge overlap between QA pairs across clinical specialties.
Finally, we ask how do clinically-relevant terms for one specialty change before and after fine-tuning the LLM? We address this question in §3.4. We again use the clinical terms curated in §3.3 and analyze changes in term probabilities before and after fine-tuning on data from the same vs. different clinical specialties. Our results suggest that performance gains are driven more by domain shifts (from general to medical domain) than by fine-grained clinical knowledge injection.
Our main contributions are:
  • We introduce S-MedQA, a medical QA dataset spanning 15 clinical specialties with 4040k QA pairs annotated with high quality and human validated specialty annotations.
  • We consider both single-specialty annotations and multiple-specialty labels to account for cross-disciplinary QA pairs.
  • We systematically evaluate the impact of fine-tuning 8 LLMs on cross-specialty medical QA performance. Our findings indicate that performance gains in medical QA tasks are primarily driven by domain shifts rather than the injection of specialty-specific knowledge.
  • We release S-MedQA and all code necessary for reproduction, providing the community with a valuable resource for benchmarking and improving medical LLMs.
Figure 2: The distribution of all specialties classified by GPT-3.5. The dark blue specialties are the 15 we finally included in our benchmark. Note that we conduct experiments on the 6 common specialties for simplicity.
2 Creation of S-MedQA
We now describe the creation of S-MedQA, a high-quality benchmark for medical QA with clinical specialty annotations. In Figure 1, we show an overview of S-MedQA’s creation process.
@@ 2 Creation of S-MedQA
Creating S-MedQA involves 1) using an LLM to produce single medical specialty annotations for each QA pair (§2.1), 2) a thorough human validation step to ascertain the quality and correctness of the annotations (§2.2), and finally 3) using an LLM to extend these single specialty annotations to possibly multiple specialties whenever the question requires so drawing on ideas from conformal prediction (§2.3). We release multiple versions of S-MedQA with varying accuracy/coverage trade-offs, controlled by majority voting thresholds for including examples. Users can opt for a cleaner dataset with fewer samples or a noisier one with more samples.
Data and splits.
@@ 2.1 Clinical specialty categorization
Manually labeling QA examples with clinical specialties is costly and time-consuming. We thus first use GPT-3.5 to annotate samples with single clinical specialties. Preliminary experiments using a single prompt to predict a single specialty produced low accuracy annotations ($\sim 75$%). To improve this, we design five prompts (details in §A.2) and predict single specialties with GPT-3.5 for each QA pair independently with each prompt. We then apply majority voting to decide on the specialty for that QA pair (Ding et al., 2023; Goel et al., 2023). Human evaluation shows accuracies between 90.890.8%–97.897.8% (more details next in §2.2).
We exclude $3,518$ (11.511.5%) samples categorized as Others since they mostly contain irrelevant clinical information (see Appendix A.4 for examples of exclusion), and then focus on $15$ out of $55$ specialties with more than $500$ samples. The final dataset comprises $24306$ / $899$ / $899$ samples in train/validation/test sets after human validation.
In Figure 2, we show the distribution of samples across specialties. We show the 15 specialties we include in S-MedQA in dark blue, comprising in total 70.0% / 70.7% / 70.1% of the entire train / valid / test sets. In Table 1, we show the 15 specialties included in S-MedQA with the respective numbers of samples. The columns represents majority voting with 3+, 4+, and 5 prompts.
Number of Votes (out of 5) 3+ 4+ 5
Cardiology 3,000 2,652 2,122
Neurology 2,954 2,419 1,627
Endocrinology 2,384 2,122 1,568
Infectious diseases 2,161 1,705 1,087
Hematology 2,005 1,746 1,332
Gastroenterology 1,934 1,594 1,176
Emergency medicine 1,597 1,098 629
Pediatrics 1,553 1,036 559
Nephrology 1,550 1,340 1,004
Respiratory medicine 1,253 825 395
Rheumatology 1,115 958 700
Obstetrics and gynecology 913 802 636
Oncology 770 570 300
Psychiatry 656 582 446
Dermatology 531 421 280
Total 24,306 19,290 14,261
Table 1: S-MedQA description. Number of samples of the 15 specialties using different minimum numbers of votes (3+, 4+, 5) in the train sets included in S-MedQA.
@@ 2.2 Manual validation for single specialty predictions
Prompts #1 #2 #3 #4 #5
Accuracy(%) 76.0 72.8 73.0 73.8 80.2
Table 2: Accuracy of each single prompt. Prompts #11 to #55 are shown in Figures 610 in Appendix A.2.
A medical expert labels all the examples in S-MedQA’s validation and test sets with the single most correct clinical specialty. This expert also validates $1,000$ random samples from the train set, confirming whether the specialties predicted by GPT-3.5 are correct. [^2] This is the specialty that is the most relevant to the QA pair. In this evaluation, we exclude any cases where there is ambiguity in which clinical specialty is the most relevant. Table 2 shows the categorization accuracies when using single prompts. The accuracies when applying majority voting are shown on top in Table 3. In general, when using voting with multiple prompts, we see large performance gains compared to using single prompts. (e.g., from 72.872.880.280.2% to 90.890.897.897.8%).
Number of Votes (out of 5) $3+$ $4+$ 55
@@
In Table 3, we also show the accuracy vs. coverage trade-off over the $1,000$ random samples from the train set for different requirements for majority voting. A higher quorum results in higher accuracy ($90.8\rightarrow 97.8$) but greatly decreases the coverage ($89.1\rightarrow 49.2$). We release per-prompt single-specialty categorizations with votes for all examples for users to decide their preference between accuracy and coverage—more data but possibly more noise or less noise but less data—based on their specific use cases. We select ‘3+’ as the quorum in this study for adequate fine-tuning data.
Figure 3: The illustration of first token probability, string matching, and our approach (classifier) to evaluating LLMs performance on S-MedQA. We use text output instead of first token probability for evaluation because the latter suffers heavily from selection bias in multiple-choice QA (Wang et al., 2024). However, string matching does not work in some cases. Our classifier trained on Mistral-v0.2 works successfully with an accuracy of 96.5%.
2.3 From single to multiple clinical specialties
We devise a strategy based on conformal prediction to expand clinical specialty annotations to possibly multiple specialties if a QA pair requires so. Our main idea is: 1) calibrate an LLM to find a threshold that maximises multiple specialties’ prediction accuracy in a calibration set; 2) use this threshold to build a conformal classifier that predicts multiple clinical specialties for each QA pair.
Calibration and test set.
We curate a held-out set of 300 samples, annotated independently by a medical expert (300 samples) and three medical students (100 samples each). We use 200 examples (66.6%) as a calibration set, and report results on the remaining 100 examples (33.3%).
Conformal threshold and set construction.
Following  Ke et al. (2025), we query GPT-3.5-Turbo $n=20$ times per input question and predict a single specialty $C_{i}$. We then get frequencies $P(C_{i})$ over all possible output specialties (see §A.1). We compute the non-conformity scores as $S_{i}=1-P(C_{i})$ and obtain a set of scores $S_{1}$, $S_{2}$, …, $S_{n}$ per question. We set our acceptable error to 10% ($\alpha=0.1$) and then use the ($1-\alpha$) quantile of all the scores as the calibration threshold. At prediction time, we also sample 20 times single specialty categorization for each question, and compute the per specialty scores. We include all specialties that have lower score than the calibration threshold in the final multi-specialty prediction.
We report key metrics computed on the testing data. The number of instances where the predicted label set exactly matched the true gold label set was $24$. Precision, defined as the proportion of predicted labels that were correct, was $0.69$. Recall, which measures the proportion of gold labels correctly included in the prediction sets, was $0.52$. The average number of predicted labels was $1.55$.
3 Experiments
We first investigate how training LLMs on data from a clinical specialty impacts their performance on other specialties (§3.2). Our findings show that training on one clinical specialty does not necessarily lead to the best performance in that specialty. We hypothesize whether QA pairs from different clinical specialties are indeed very different from each other. That leads us to quantify the overlap of clinically relevant terms in train/test splits within and across specialties (§3.3). Finally, we investigate to what extent token probabilities of clinically relevant tokens change when fine-tuning LLMs in S-MedQA compared to non-clinical data (§3.4).
@@ 3 Experiments
Train Sets Llama-3.1 64.0 67.5 65.3 66.8 65.0 64.5 65.5
Cardio 69.3 76.5 73.5 76.2 70.5 70.9 72.8
Gastro 68.2 73.3 71.8 74.6 69.3 69.2 71.1
Infect 65.8 74.2 69.5 72.5 66.6 66.7 69.3
Neuro 69.9 76.4 72.3 76.9 69.5 71.7 72.8
Obstetrics 69.8 75.0 72.9 75.4 68.7 69.4 71.9
Pediatrics 66.9 75.0 72.1 75.9 68.7 70.2 71.4
Combined 72.9 79.5 75.4 77.8 70.8 74.1 75.1
Table 4: Accuracy matrix for Llama-3.1-8B-Instruct. Model is applied without finetuning. Model is trained on the combination of all 6 specialty train sets. For robustness, accuracies are computed as the average over the test set with answer options shuffled five times. For each specialty, the best performance when fine-tuned on different specialty datasets is in bold, and scores for models fine-tuned on the same specialty are underlined. Surprisingly, none of the best performances come from models fine-tuned on their corresponding training sets.
@@ 3.1 Experimental setup
To minimize potential evaluation bias, we follow best practices to evaluate LLM performance on QA datasets (1) by shuffling the order of the answers multiple times and adding multiple shuffled QA pairs in the test set, and (2) by using the entire answer to a question instead of the LLM’s maximum token probability among options A, B, C, D in the answer (Zheng et al., 2023; Wang et al., 2024).
More specifically, in all test sets, we shuffle the answers 5 times for each sample and add all these 5 entries to the final test set in case the model prefers an option due to its position (Zheng et al., 2023). To further improve the reliability, we follow Wang et al. (2024) to train a classifier to match model outputs to the options in a post-hoc step, instead of using the maximum probability of options {A, B, C, D} with a single next-token prediction step. More concretely, we randomly select 150 training samples and generate answers for these with four LLMs (Llama2-7b, Llama2-13b, Mistral-v0.1, and Mistral-v0.2), resulting in $600$ responses. We manually annotate all the responses with the right options and use these annotations to train a Mistral-Instruct-v0.2 model as the classifier, with $400$ ($200$) train (test) samples. Our classifier achieves 96.5% accuracy and we use it in all experiments.
In all subsequent analyses below, we use the 6 most common specialties (Cardiology, Gastroenterology, Infectious diseases, Neurology, Obstetrics and Gynecology, and Pediatrics), with test set sample size of $80$, $83$, $102$, $74$, $88$, and $90$, respectively.
3.2 Cross-specialty evaluation
We experiment with 8 open-source LLMs: Llama2-Chat-7B, Llama2-Chat-13B, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct (Touvron et al., 2023), Mistral-Instruct-v0.1, Mistral-Instruct-v0.2 (Jiang et al., 2023) , Bio-Medical-LLaMA-3-8B (Con, 2024), and OLMo-2-1124-7B-Instruct (OLMo et al., 2024). These include six general-purpose models (Llama-2, Llama-3, and Mistral-based models), one biomedical Llama model, and a fully open-source model (OLMo).
We fine-tune each LLM on the six per-specialty training sets using prompts from §A.3, and evaluate on all six test sets. Models are trained for up to 10 epochs, with selection based on per-specialty validation accuracy. We also train on the combined dataset to evaluate how exposure to larger and more diverse data affects models’ performance. We use LoRA (Hu et al., 2021) on all projection layers for the fine-tuning process in all experiments with the following hyperparameters: learning rate 22e-55, rank $32$, alpha $16$, dropout rate $0.1$, batch size $8$.
Results.